# State-of-the-Art Dysarthric Speech Recognition with MetaICL for on-the-fly Personalization

### v1

@@ 4 Conclusion
This paper presents a robust and scalable solution for on-the-fly personalization of dysarthric speech recognition. Our key contribution is a novel mixed-objective meta-learning strategy that fine-tunes a multi-modal LLM to excel at both zero-shot transcription and few-shot, in-context personalization. We empirically prove the superiority of this hybrid approach over single-objective methods and establish new state-of-the-art benchmarks on the diverse Euphonia and SAP datasets, surpassing prior work that relies on complex, per-user personalized models. Our investigation into enrollment curation—showing that 5 examples selected with an oracle method can match the performance of 19 random ones—underscores the potential for high efficiency. This finding strongly motivates future work in developing deployable curation methods, such as those based on acoustic similarity, to make on-the-fly personalization even more practical. Lastly, our findings from the data ablation study show an efficient, front-loaded learning curve where just 2% of training data delivered over 70% of the performance gains. By eliminating the need for separate model training and storage, this work offers a clear and effective path toward more accessible and easily deployable ASR systems that can adapt to any user instantly.
5 Acknowledgments
The authors would like to acknowledge the providers of the Euphonia dataset (Google Research and collaborators) and the Speech Accessibility Project dataset (led by UIUC) for making such valuable resources available to the research community; and thank Jimmy Tobin, Subhashini Venugopalan, Rohit Prabhavalkar, Dana Alon, and Kurt Partridge for the helpful discussions.

### v2

@@ 4 Conclusion
This paper presents a robust and scalable solution for on-the-fly personalization of dysarthric speech recognition. Our key contribution is a mixed-objective meta-learning strategy that fine-tunes a multi-modal LLM to excel at both zero-shot transcription and few-shot, in-context personalization. We establish new state-of-the-art benchmarks on Euphonia and SAP Test-1, and achieve competitive performance on SAP Test-2, all without requiring user-specific weight updates, offline model-merging, or custom audio segmentation.
Our analysis of enrollment curation validates the model’s active adaptation, evidenced by a 40% WER reduction when using random same-speaker support examples. We find this random selection forms a surprisingly robust baseline that static, text-based curation strategies fail to beat. However, our oracle experiments reveal efficiency headroom when examples are matched via similarity, highlighting dynamic acoustic retrieval as the next critical frontier for this technology.
Lastly, our findings from the data ablation study show that the model learns to utilize in-context examples with only 1000 total instances, and domain adaptation although slower, also takes only 20,000 mixed objective instances to reach a desirable level. This work opens the door for instant, high-quality ASR for the millions of individuals living with speech impairments, without the burden of a long and exhausting enrollment process.